1.      A. Define elasticity of supply and find the price from the given statement:

If ES of a good 2and a firm supplies 200 units at price of Rs 8 per unit, then at what price will the firm supply 250 units.

B. Calculate the elasticity of supply if a 15% increase in the price of soya bean oil increases its supply from 300 to 345 units. 


1
Expert's answer
2021-05-14T10:55:28-0400

A. Price elasticity of supply measures the responsiveness to the supply of a good or service after a change in its market price. Elastic means the product is considered sensitive to price changes. Inelastic means the product is not sensitive to price movements.

"Ed = \frac{250 - 200} {P2 - 8} ×\frac{P2 + 8} {250+200} = 2,"

18P2 - 144 = P2 + 8,

17P2 = 152,

P2 = 8.94.

B. The increases in supply is 345/300 - 1 = 0.15 or 15% and is the same as an increase in price.

So, Ed = 1.